If you could help me out with this problem that would be amazing! thanks for any helpJorge and Anita, married taxpayers, earn $140,500 in taxable income and $46,000 in interest from an investment n Cty of Heflin bonds. (Use the U.S. tax rate schedule for married filing jointly) Required a. If Jorge and Anita earn an additional $103,000 of taxable income, what is their marginal tax rate on this income? b. What is their marginal rate if, instead, they report an additional $103,000 in deductions? (For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places.) a. Marginal tax rate b Marginal tax rate

Schedule Y-1-Married Filing Jointly or Qualitying Widow(er) If taxable income is over: But not over: 19,050 $ 77,400 $165,000 $315,000 $400,000 $600,000 The tax is: S 19,050 S 77,400 $165.000 $315,000 $400,000 $600,000 10% of taxable income $1,905 plus 12% of the excess over $19,050 $3.907 plus 22% of the excess over $77,400 $28,179 plus 24% of the excess over $165,000 S64-179 plus 32% of the excess over $315,000 $91,379 plus 35% of the excess over $400,000 $161,379 plus 37% of the excess over $600,000
Answer #1
Tax on $140500 = 8907+(140500-77400)*22%= $22789
a
Taxable income = 140500+103000= $243500
Tax on $243500 = 28179+(243500-165000)*24%= $47019
Marginal tax rate = (47019-22789)/103000= 23.52%
b
Taxable income = 140500-103000= $37500
Tax on $37500 = 1905+(37500-19050)*12%= $4119
Marginal tax rate = (22789-4119)/103000= 18.13%